OpenStudy (anonymous):

The strength of an electric field at 5.0 cm from a point charge is 100.0 N/C. What is the magnitude of the source charge? Show your work.

OpenStudy (anonymous):

The electric field is given by the following equation:\[E = \frac{1}{4 \pi \epsilon_0} \frac{q}{r} = 100 N/C\] where r = 0.05 m, \[\epsilon_0 = 8.85 \times 10^{-12}\] and q is the source charge. So by replacing these values into the equation, we find \[q = 5.56 \times 10^{-10} C\].